# Dynamics of Moral Behavior in Heterogeneous Populations of Learning Agents

### v3

@@ Abstract
Growing concerns about safety and alignment of AI systems highlight the importance of embedding moral capabilities in artificial agents. A promising solution is the use of learning from experience, i.e., Reinforcement Learning. In multi-agent (social) environments, complex population-level phenomena may emerge from interactions between individual learning agents. Many of the existing studies rely on simulated social dilemma environments to study the interactions of independent learning agents. However, they tend to ignore the moral heterogeneity that is likely to be present in societies of agents in practice. For example, at different points in time a single learning agent may face opponents who are consequentialist (i.e., caring about maximizing some outcome over time) or norm-based (i.e., focusing on conforming to a specific norm here and now). The extent to which agents’ co-development may be impacted by such moral heterogeneity in populations is not well understood. In this paper, we present a study of the learning dynamics of morally heterogeneous populations interacting in a social dilemma setting. Using a Prisoner’s Dilemma environment with a partner selection mechanism, we investigate the extent to which the prevalence of diverse moral agents in populations affects individual agents’ learning behaviors and emergent population-level outcomes. We observe several types of non-trivial interactions between pro-social and anti-social agents, and find that certain classes of moral agents are able to steer selfish agents towards more cooperative behavior.
1 Introduction
@@
In general, it has been shown that morality can be developed in agents through learning from experience (Tennant et al., 2023b). Moral principles can range from consequentialist (Bentham, 1780) to norm-based morality (Kant, 1785), and from entirely pro-social to entirely anti-social preferences. Existing works have proposed a variety of individual models to predispose agents to specific moral principles (e.g., Bazzan et al. 1999; Capraro & Perc 2021), and recent studies have demonstrated that these can be effectively encoded in Reinforcement Learning (RL) agents via intrinsic rewards (Hughes et al., 2018; McKee et al., 2020; Tennant et al., 2023a).
@@
However, to date we still have limited understanding of how different types of morality may co-evolve in heterogeneous populations. In fact, many real-world AI systems are likely to co-exist (essentially forming systems of systems) and may be co-developed in parallel with others. Especially when it comes to integrating moral decision-making into systems, different stakeholders may decide to prioritize varying principles or preferences. It is therefore crucial that we start developing an understanding of how the presence of different moral preferences in populations affects individual agents’ learning behaviors and interactions, and, indeed, emergent population-level outcomes.
@@
This type of agent-based analysis of moral learning may additionally offer insights into the potential dynamics of interactions in human societies in the tradition of computational philosophy (McKenzie, 2007), provide an experimental test-bed for new theories (Mayo-Wilson & Zollman, 2021), or simulate potential evolutionary mechanisms underpinning human moral preferences, in a similar way to Evolutionary Game Theory (Hofbauer & Sigmund, 1998; Sigmund & Nowak, 1999; Sigmund, 2010).
@@
The core contribution of this work is the study of the behavior and population dynamics among RL agents with diverse moral preferences, providing insights for artificial agent design also in terms of safety [^1] Code: will be made available upon acceptance. More generally, we present a methodology for analyzing emergent behavior in populations of agents with heterogeneous moral orientation.
2 Background
@@
Table 1: Payoffs on one step of the Iterated Prisoner’s Dilemma (IPD).
2.1 Social Dilemma Games
Social dilemma games simulate social situations in which players obtain different utilities (payoffs) from choosing one action or another, and the structure of these utilities is such that each player faces a trade-off between individual interest and societal benefit when choosing an action (Dawes, 1980). The most widely studied type is a symmetric matrix game with two players and two abstract actions - Cooperate or Defect. Players in these games must decide on their respective actions simultaneously, without communicating.
A classic game from Economics and Philosophy, which is relevant to moral choice, is the Prisoner’s Dilemma (Rapoport 1974, see payoffs for row vs column player in Table 1). We implement the Iterated Prisoner’s Dilemma (IPD) in a population, in which agents interact in pairs in discrete time steps and aim to maximize their cumulative payoff over time. In the iterated game, players can take actions to punish their opponents for past defection or to influence their future behaviors.
@@ 2.1 Social Dilemma Games
Deriving predicted equilibria for these games is not always feasible; indeed, since these are systems composed of interacting (heterogeneous) entities, simulation is needed to study potential emergent behaviors and outcomes (Anderson, 1972). Additionally, learning dynamics may cause instability in the environment and dynamic behaviors (Busoniu et al., 2008).
2.2 Deep Reinforcement Learning in Markov Games
@@
Reinforcement Learning (RL) is a well-suited technique for modeling agents that learn by interacting with others in an environment (Sutton & Barto, 2018). It can be applied in conjunction with Evolutionary Game Theory (Hofbauer & Sigmund, 1998) to iterated social dilemma games (Littman, 1994; Sandholm & Crites, 1996; de Cote et al., 2006; Abel et al., 2016), in which game payoffs constitute extrinsic rewards, and traits such as moral or social preferences (Fehr & Fischbacher, 2002) can be encoded in the agent’s intrinsic reward (Chentanez et al., 2004). In the following section, we discuss the design of intrinsic moral rewards in detail. It is also worth noting that we assume populations of independent, continuously learning agents. This creates interesting dynamics as agents affect one another’s learning process (Leibo et al., 2019). Given the potentially large number of states, we adopt DQN as the underlying learning algorithm (Mnih et al., 2015).
2.3 Morality as Intrinsic Reward
@@
Traditional social dilemma scenarios assume that agents are only motivated by accumulating game payoffs (i.e., extrinsic reward). However, human data has shown that other preferences may also come into play, such as the predisposition to cooperation (Camerer, 2011). In artificial agents, preferences other than game rewards can be encoded in intrinsic rewards (Chentanez et al., 2004). As a result, recent multi-agent RL studies have become increasingly interested in modeling various pro-social preferences as intrinsic reward functions for social dilemma players (Hughes et al., 2018; Peysakhovich & Lerer, 2018; McKee et al., 2020). Moral rewards in particular have been studied for two-agent (i.e., dyadic) interactions in Tennant et al. (2023a).
@@
We anchor our intrinsic reward definitions in traditional moral philosophical frameworks, especially relying on the distinction between consequentialist versus norm-based morality, and the idea of virtue ethics. Consequentialist morality focuses on the consequences of an action, and includes Utilitarianism (Bentham, 1780), which defines actions as moral if they maximize total utility for all agents in a society. Norm-based morality, including Deontological ethics (Kant, 1785), considers an act moral if it does not contradict the society’s external norms. Finally, in Virtue Ethics (Aristotle, ), moral agents must act in line with their certain internal virtues, such as fairness or care for others (Graham et al., 2013). Different virtues can matter more or less to different agents (Aristotle, ; Graham et al., 2009) and can themselves have consequentialist or norm-based foundations. We present the set of agents considered in this study and their classification in Section 3.2.
2.4 Partner Selection
@@
In human populations, agents have a choice of which individual to interact with. Given this selection mechanism, reputation comes into play and competitive and collaborative relationships may form. Santos et al. (2008) show experimentally that these dynamics are likely to lead to a re-structuring of the population. Adaptive behavior resulting from selection mechanisms has been hypothesized to drive the emergence of cooperation (Barclay & Willer, 2007; Cuesta et al., 2015). Partner selection mechanisms and emergence of cooperation have recently been studied in societies of artificial learning agents as well (Anastassacos et al., 2020; Baker, 2020).
@@
In particular, applying partner selection to the IPD, a conflict may arise between a player’s motivation to appear cooperative in order to get selected more often by other payers, and the motivation to select and then exploit cooperators to gain a greater payoff. Anastassacos et al. (2020) find partner selection to be norm-inducing and even lead to the emergence of cooperation in a population of purely selfish agents. We adopt a partner selection model similar to Anastassacos et al. (2020).
@@
Selection dynamics in morally heterogeneous populations are likely to be more complex and harder to predict. Since each player plays according to their own intrinsic reward signal, different coalitions (Shenoy, 1979) may arise among different agent types within a population, and popularity of agents may not directly correlate with cooperativeness. Running simulation experiments to study this may provide insight into the types of behaviors and outcomes that may develop in heterogeneous populations with selection, with implications for the study of human and animal societies and Evolutionary Game Theory (Sigmund & Nowak, 1999).
2.5 Learning in Heterogeneous Populations
@@
Evidence from the social sciences suggests that human societies are morally heterogeneous (Graham et al., 2013; Bentahila et al., 2021). While certain commonsense norms may be agreed upon by a society (Reid & Haakonssen, 1990; Gert, 2004), distinct moral principles are likely to hold a different weight for different individuals depending on factors such as political orientation (Graham et al., 2009), personality (Lifton, 1985), nurture or even natural predispositions (Sinnott-Armstrong, 2008).
@@
Another social dilemma study with diverse consequentialist agents (i.e., agents with various preferences over group outcome distributions, McKee et al. 2020) suggests that agents trained in heterogeneous populations develop particularly generalized and high-performing policies. In contrast, populations containing only altruistic agents are characterized by greater collective reward, but at the cost of equality between agents. In McKee et al. (2022), the authors define population diversity as the set of various opponents’ policies that an agent may face, and also find that training in diverse populations can lead to improvements in terms of agent performance on some types of environments.
@@
Even without learning, Santos et al. (2008) show that - in fixed networks of agents - diversity in terms of neighborhoods within the population graph promotes cooperation in the population. With respect to global societal outcomes, Ord (2015) discusses how, through interaction, morally diverse agents taking actions to satisfy their distinct moral goals may improve the global welfare of the whole population, akin to a ‘trade’. We believe that these conceptual frameworks can represent the basis for studying how artificial agents may learn to act across populations containing different proportions of opponents with different moral preferences.
3 Methodology
3.1 Experimental Setup
@@
Our environment involves a population $P$ (of size $N=16$) playing an iterated game. At each iteration, a moral player $M$ and an opponent $O$ play a one-shot Prisoner’s Dilemma game with two possible actions - Cooperate or Defect ($a^{t}_{M},a^{t}_{O}\in\{C,D\}$, see Table 1 for the respective payoffs). To allow learning in a population, we run our simulation in 30000 episodes, where a single episode involves each player selecting a partner once, and then each corresponding pair of players playing the game. In particular, at iteration $t$, the learning agent $M$ observes a selection state including the action played by each possible opponent $O\in P$ at $t-1$: $s_{M,sel}^{t}=[a_{O_{1}}^{t-1},...,a_{O_{n}}^{t-1}]$. Using this state and a policy, $M$ selects an opponent $O$. Then, to choose an action for the dilemma, player $M$ relies on the state representing the latest move of their selected opponent $O$: $s_{M,dil}^{t}=a_{O}^{t-1}$. Simultaneously, their opponent $O$ chooses an action following the same principle. The players $M$ and $O$ then each receive a game reward $R_{M}^{t+1}$ and $R_{O}^{t+1}$ (corresponding to the game payoffs) and observe a new state $s_{M,dil}^{t+1}$ and $s_{O,dil}^{t+1}$ based on their opponent’s move $a^{t}$.
@@
In this Markov game (Littman, 1994), over time both players learn to make selections and take actions by observing past interactions with various opponents. Each agent uses Q-Learning (Watkins & Dayan, 1992) to update the state-action value function $Q$ (i.e., an estimate of the cumulative expected return over time). Given a large number of possible selection states, we approximate $Q$ using a neural network; each agent maintains two internal models (one for partner selection, and one for playing the dilemma) - for each of these we use a fully connected network with a single hidden layer of size $265$.
For learning both selections and the dilemma, players record the experiences $(s,a,r,s^{\prime})$ in their memory buffer $D_{sel}$ and $D_{dil}$. We copy the reward obtained from the game into both the dilemma and selection memories. At the end of an episode, all players simultaneously update the Q-network parameters $\theta_{t}$ for their two models using the latest experience available in the memory buffer $D$, using the Mean Squared Error loss function (Mnih et al., 2015) where $\gamma=0.99$ is a discount factor:
$L_{t}(\theta_{t})=\mathbb{E}_{s,a,r,s^{\prime}}\bigg{[}\bigg{(}R^{t+1}+\gamma\max_{a}Q(s^{t+1},a,\theta_{t})-Q(s^{t},a^{t},\theta_{t})\bigg{)}^{2}\bigg{]}$ (1)
If more than one experience is available (i.e., a player plays more than one dilemma game in that episode), we calculate an average loss across the experiences before updating the network. Each memory buffer $D$ is refreshed before the start of the next episode, so that each player only learns from the latest episode of experience. In updating the network weights with the Adam optimizer (Kingma & Ba, 2015), we use a learning rate of $0.001$.
Agents select partners and play the dilemma using an $\epsilon$-greedy policy, acting randomly with probability $\epsilon$ and otherwise playing greedily according to the Q-values learned so far:
We set $\epsilon_{sel}=0.1$ and $\epsilon_{dil}=0.05$. The full detailed algorithm can be found in Appendix A.
Player $M$ can learn according to an extrinsic game reward $R_{M_{extr}}^{t+1}$, which depends directly on the joint actions $a_{M}^{t},a_{O}^{t}$ (as defined in Table 1), or according to an intrinsic reward $R_{M_{intr}}^{t+1}$. In the next subsection, we discuss the definition of these intrinsic moral rewards.
3.2 Modeling Morality as Intrinsic Reward
@@
Ut Utilitarian $R_{M_{intr}}^{t}=R_{M_{extr}}^{t}+R_{O_{extr}}^{t}$ External/Internal conseq.
@@
aUt Anti-Utilitarian $R_{M_{intr}}^{t}=-(R_{M_{extr}}^{t}+R_{O_{extr}}^{t})$ External/Internal conseq.
@@
Table 2: Definitions of the types of intrinsic moral rewards, from the point of view of the moral agent $M$ playing versus an opponent $O$ at iteration $t$. We define four pro-socially-disposed players, four anti-social ones, and the traditional Selfish player.
@@
Intrinsic rewards associated with pro-social preferences have been used to incentivize the emergence of cooperation in social dilemmas (Hughes et al., 2018; Peysakhovich & Lerer, 2018; Jaques et al., 2019). Tennant et al. (2023a) extended this work by defining four Q-learning moral [^2] We use the term moral for agents that are not selfish. However, selfishness itself can be considered as a moral choice, expressed as rational egotism. In the case of social dilemmas, selfishness maps to the concept of rationality. For these agents $R_{M_{intr}}=R_{M_{extr}}$. agents that learn according to various intrinsic rewards $R_{M_{intr}}$, and contrasting these against a traditional Selfish agent which learns to maximize its extrinsic (game) reward $R_{M_{extr}}$ (e.g., Leibo et al. 2017). The formal definitions of the reward functions are presented in Table 2.
@@
In this study, we rely on eight types of moral agents - four pro-social moral learners from Tennant et al. (2023a) (Utilitarian, Deontological, Virtue-Equality, Virtue-Kindness), and four additional anti-social counterparts (anti-Utilitarian, malicious-Deontological, Virtue-Inequality, Virtue-Aggression):
@@
  • the Utilitarian agent tries to maximize the collective payoff (i.e., total payoff for both players; Bentham 1780);
@@
  • the Deontological agent tries to follow the conditional cooperation norm (Kant, 1785; Fehr & Fischbacher, 2004) and gets penalized through the negative reward $-\xi$ for defecting against a cooperator (i.e., an opponent who previously cooperated);
@@
  • the Virtue-Equality agent tries to maximize equality between the two players’ payoffs $R_{M_{extr}}^{t}$ and $R_{O_{extr}}^{t}$, measured using a two-agent variation of the Gini coefficient (Gini, 1912);
@@
  • the Virtue-Kindness agent receives a reward $\xi$ for acting kindly (i.e., cooperating) against any opponent (Aristotle, );
  • the anti-Utilitarian agent tries to minimize the collective payoff;
@@
  • the malicious-Deontological agent tries to follow the conditional defection norm and receives a reward $\xi$ for defecting against an cooperator;
@@
  • the Virtue-Inequality agent tries to minimize equality between the two payoffs;
@@
  • the Virtue-Aggression agent receives a reward $\xi$ for acting aggressively (i.e., defecting) against any opponent.
@@ 3.3 Population Compositions
Table 3: Populations considered in this study. Each population contains eight players of a ‘majority’ type and one player of each other type.
4 Results

### v4

@@ Abstract
Growing concerns about safety and alignment of AI systems highlight the importance of embedding moral capabilities in artificial agents: a promising solution is the use of learning from experience, i.e., Reinforcement Learning. In multi-agent (social) environments, complex population-level phenomena may emerge from interactions between individual learning agents. Many of the existing studies rely on simulated social dilemma environments to study the interactions of independent learning agents; however, they tend to ignore the moral heterogeneity that is likely to be present in societies of agents in practice. For example, at different points in time a single learning agent may face opponents who are consequentialist (i.e., focused on maximizing outcomes over time), norm-based (i.e., conforming to specific norms), or virtue-based (i.e., considering a combination of different virtues). The extent to which agents’ co-development may be impacted by such moral heterogeneity in populations is not well understood. In this paper, we present a study of the learning dynamics of morally heterogeneous populations interacting in a social dilemma setting. Using an Iterated Prisoner’s Dilemma environment with a partner selection mechanism, we investigate the extent to which the prevalence of diverse moral agents in populations affects individual agents’ learning behaviors and emergent population-level outcomes. We observe several types of non-trivial interactions between pro-social and anti-social agents, and find that certain types of moral agents are able to steer selfish agents towards more cooperative behavior.
Introduction
@@
In general, it has been shown that morality can be developed in agents through learning from experience (Hughes et al. 2018; McKee et al. 2020; Tennant, Hailes, and Musolesi 2023b). Learning approaches to moral alignment offer a variety of advantages, including potentially greater adaptability and generality (Tennant, Hailes, and Musolesi 2023a), and the ability to learn implicit preferences (Ouyang et al. 2022). Moral principles can range from consequentialist (Bentham 1780) to norm-based morality (Kant 1785), and from entirely pro-social to entirely anti-social preferences. Furthermore, the morality of agents can be based on a combination of virtues, as exemplified by the Virtue Ethics approach (Aristotle n.a.). Existing works have proposed a variety of individual models to predispose agents to specific moral principles (e.g., Bazzan, Bordini, and Campbell 1999; Capraro and Perc 2021), and recent studies have demonstrated that these can be effectively encoded in Reinforcement Learning (RL) agents via intrinsic rewards (Hughes et al. 2018; McKee et al. 2020; Tennant, Hailes, and Musolesi 2023b). Version with Appendix: https://arxiv.org/abs/2403.04202 Code: https://github.com/Liza-Tennant/moral˙MARL
@@
However, to date, we still have limited understanding of how different types of morality may co-evolve in heterogeneous populations. In fact, many real-world AI systems are likely to co-exist (essentially forming systems of systems) and may be co-developed in parallel with others. Especially when it comes to integrating moral decision-making into systems, different stakeholders may decide to prioritize varying principles or preferences. It is therefore crucial that we start developing an understanding of how the presence of different moral preferences in populations affects individual agents’ learning behaviors and interactions, and, indeed, emergent population-level outcomes. Furthermore, simulation studies such as ours contribute to raising awareness about emergent behaviors and the possibility of unintuitive outcomes emerging from multi-agent learning scenarios.
@@
This type of agent-based analysis of moral learning may additionally offer insights into the potential dynamics of interactions in human societies in the tradition of computational philosophy (McKenzie 2007), provide an experimental test-bed for new theories (Mayo-Wilson and Zollman 2021), or simulate potential evolutionary mechanisms underpinning human moral preferences, in a similar way to Evolutionary Game Theory (Hofbauer and Sigmund 1998; Sigmund and Nowak 1999; Sigmund 2010).
@@
The core contribution of this work is the study of behavior and population dynamics among RL agents with diverse moral preferences, providing insights for the design of artificial agents with a focus on safety and alignment. More generally, we present a methodology for analyzing emergent behavior in populations of agents with heterogeneous moral orientation. We apply the proposed methodology to the Iterated Prisoner’s Dilemma (IPD, Rapoport 1974), demonstrating at the same time its potential applicability and generalizability to a variety of scenarios and different moral frameworks.
Background and Preliminaries
Social Dilemma Games
Social dilemma games simulate social situations in which players obtain different utilities (payoffs) from choosing one action or another, and the structure of these utilities is such that each player faces a trade-off between individual interest and societal benefit when choosing an action (Dawes 1980). The most widely studied type is a symmetric matrix game with two players and two abstract actions - Cooperate (C) or Defect (D). Players in these games must decide on their respective actions simultaneously, without communicating.
A classic game from Economics and Philosophy, which is relevant to moral choice, is the Prisoner’s Dilemma (Rapoport 1974, see payoffs for row vs column player in Table 1). We implement the Iterated Prisoner’s Dilemma (IPD) in a population, in which agents interact in pairs in discrete time steps and aim to maximize their cumulative payoff over time. In the iterated version of the game, players can take actions to punish their opponents for past defection or to influence their future behaviors.
@@
Table 1: Payoffs on one step of the Iterated Prisoner’s Dilemma (IPD) game for the row & column player.
@@
Deriving predicted equilibria for these games is not always feasible: these are systems composed of interacting (heterogeneous) entities, where learning dynamics may cause instability in the environment and dynamic behaviors (Busoniu, Babuska, and De Schutter 2008). Therefore, simulation is needed to study potential emergent behaviors and outcomes (Anderson 1972).
Deep Reinforcement Learning in Markov Games
@@
Reinforcement Learning (RL) is a well-suited technique for modeling agents that learn by interacting with others in an environment (Sutton and Barto 2018). It can be applied in conjunction with Evolutionary Game Theory (Hofbauer and Sigmund 1998) to iterated social dilemma games (Littman 1994; Sandholm and Crites 1996; de Cote, Lazaric, and Restelli 2006; Abel, MacGlashan, and Littman 2016), in which game payoffs constitute extrinsic rewards, and traits such as moral or social preferences (Fehr and Fischbacher 2002) can be encoded in the agent’s intrinsic reward (Chentanez, Barto, and Singh 2004). In the following section, we discuss the design of intrinsic moral rewards in detail. It is also worth noting that we assume populations of independent, continuously learning agents. This creates interesting dynamics as agents affect one another’s learning process (Leibo et al. 2019). Given the potentially large number of states, we adopt DQN as the underlying learning algorithm (Mnih et al. 2015).
Morality as Intrinsic Reward
@@
Traditional social dilemma scenarios assume that agents are only motivated by accumulating game payoffs (i.e., extrinsic reward). However, human data has shown that other preferences may also come into play, such as the predisposition to cooperation (Camerer 2011). In artificial agents, preferences other than game rewards can be encoded in intrinsic rewards (Chentanez, Barto, and Singh 2004). As a result, recent multi-agent RL studies have been focused on modeling various pro-social preferences as intrinsic reward functions for social dilemma players (Hughes et al. 2018; Peysakhovich and Lerer 2018; McKee et al. 2020). Moral rewards in particular have been studied for two-agent (i.e., dyadic) interactions in Tennant, Hailes, and Musolesi (2023b).
@@
We anchor our intrinsic reward definitions in traditional moral philosophical frameworks, especially relying on the distinction between consequentialist versus norm-based morality, and the idea of virtue ethics. Consequentialist morality focuses on the consequences of an action, and includes Utilitarianism (Bentham 1780), which defines actions as moral if they maximize total utility for all agents in a society. Norm-based morality, including Deontological ethics (Kant 1785), considers an act moral if it does not contradict the society’s external norms. Finally, in Virtue Ethics (Aristotle n.a.), moral agents must act in line with their certain internal virtues, such as fairness or care for others (Graham et al. 2013). Different virtues can matter more or less to different agents (Aristotle n.a.; Graham, Haidt, and Nosek 2009) and can themselves have consequentialist or norm-based foundations. We present the specific set of agents considered in this study and their classification in the Methodology section.
Partner Selection
@@
In human populations, agents have a choice of which individual to interact with. Given this selection mechanism, reputation comes into play and competitive and collaborative relationships may form. Santos, Santos, and Pacheco (2008) show experimentally that these dynamics are likely to lead to a re-structuring of the population. Adaptive behavior resulting from selection mechanisms has been hypothesized to drive the emergence of cooperation (Barclay and Willer 2007; Cuesta et al. 2015). In the context of societies of artificial learning agents, these mechanisms have recently been studied by Anastassacos, Hailes, and Musolesi (2020) and Baker (2020).
@@
In particular, when applying partner selection to the IPD, a conflict may arise between a player’s motivation to appear cooperative in order to get selected more often by other payers, and the motivation to select and then exploit cooperators to gain a greater payoff. Anastassacos, Hailes, and Musolesi (2020) find partner selection to be norm-inducing and even lead to the emergence of cooperation in a population of purely selfish agents. We adopt a partner selection model similar to Anastassacos, Hailes, and Musolesi (2020).
@@
Selection dynamics in morally heterogeneous populations are likely to be more complex and harder to predict. Since each player plays according to their own intrinsic reward signal, different coalitions (Shenoy 1979) may arise among different agent types within a population, and popularity of agents may not directly correlate with cooperativeness. Running simulation experiments to study this may provide insight into the types of behaviors and outcomes that may develop in heterogeneous populations with selection, with implications not only for the study of artificial agents, but also of human and animal societies and Evolutionary Game Theory (Sigmund and Nowak 1999).
Learning in Heterogeneous Populations
@@
Evidence from the social sciences suggests that human societies are morally heterogeneous (Graham et al. 2013; Bentahila, Fontaine, and Pennequin 2021). While certain commonsense norms may be agreed upon by a society (Reid and Haakonssen 1990; Gert 2004), distinct moral principles are likely to hold a different weight for different individuals depending on factors such as political orientation (Graham, Haidt, and Nosek 2009), personality (Lifton 1985), nurture or even natural predispositions (Sinnott-Armstrong 2008).
@@
McKee et al. (2020) conducted a social dilemma study with diverse consequentialist agents (i.e., agents with various preferences over group outcome distributions). Their findings suggest that agents trained in heterogeneous populations develop particularly generalized and high-performing policies. In contrast, populations containing only altruistic agents are characterized by greater collective reward, but at the cost of equality between agents. In addition, in McKee et al. (2022), the authors define population diversity as the set of various opponents’ policies that an agent may face, and also find that training in diverse populations can lead to improvements in terms of agent performance on some types of environments.
@@
Even without learning, Santos, Santos, and Pacheco (2008) show that - in fixed networks of agents - diversity in terms of neighborhoods within the population graph promotes cooperation in the population. With respect to global societal outcomes, Ord (2015) discusses how, through interaction, morally diverse agents taking actions to satisfy their distinct moral goals may improve the global welfare of the whole population, akin to a ‘trade’. We believe that these conceptual frameworks can represent the basis for studying how artificial agents may learn to act across populations containing different proportions of opponents with different moral preferences.
Methodology
Experimental Setup
@@
Our environment involves a population $P$ (of size $N=16$) playing an iterated game. At each iteration, a moral player $M$ and an opponent $O$ play a simultaneous one-shot Prisoner’s Dilemma game with two possible actions - Cooperate or Defect ($a^{t}_{M},a^{t}_{O}\in\{C,D\}$, see Table 1 for the respective payoffs). To allow learning in a population, we run our simulation in 30000 episodes, where a single episode involves each player selecting a partner once, and then each corresponding pair of players playing the game. For clarity, each agent gets to make a selection on each episode, so $N$ games are played. In particular, at iteration $t$, the learning agent $M$ observes a selection state including the action played by each possible opponent $O\in P$ at $t-1$: $s_{M,sel}^{t}=[a_{O_{1}}^{t-1},...,a_{O_{n-1}}^{t-1}]$. Using this state and the current learned policy, $M$ selects an opponent $O$. Then, to choose an action for the dilemma, player $M$ relies on the state representing the latest move of their selected opponent $O$: $s_{M,dil}^{t}=a_{O}^{t-1}$. Simultaneously, their opponent $O$ chooses an action following the same principle. The players $M$ and $O$ then each receive a game reward $R_{M}^{t+1}$ and $R_{O}^{t+1}$ (corresponding to the game payoffs) and observe a new state $s_{M,dil}^{t+1}$ and $s_{O,dil}^{t+1}$ based on their opponent’s move $a^{t}$.
@@
In this Markov game (Littman 1994), over time both players learn to make selections and take actions by observing past interactions with various opponents. Each agent uses Q-Learning (Watkins and Dayan 1992) to update the state-action value function $Q$ (i.e., an estimate of the cumulative expected return over time). Given a large number of possible selection states, we approximate $Q$ using a neural network. Each agent maintains two internal models (one for partner selection, and one for playing the dilemma) - for consistency, we use function approximation for learning both models, and each is parametrized by a fully connected network with a single hidden layer of size $256$.
@@
Ut Utilitarian $R_{M_{intr}}^{t}=R_{M_{extr}}^{t}+R_{O_{extr}}^{t}$ External/Internal consequentialist
@@
aUt Anti-Utilitarian $R_{M_{intr}}^{t}=-(R_{M_{extr}}^{t}+R_{O_{extr}}^{t})$ External/Internal consequentialist
@@
Table 2: Definitions of the types of intrinsic moral rewards, from the point of view of the moral agent $M$ playing versus an opponent $O$ at iteration $t$. We define four pro-social players, four anti-social ones, and the traditional Selfish player.
For learning both selections and the dilemma, players record the experiences $(s,a,r,s^{\prime})$ in their memory buffer $D_{sel}$ and $D_{dil}$. We copy the reward obtained from the game into both the dilemma and selection memories. At the end of an episode, all players simultaneously update the Q-network parameters $\theta_{t}$ for their two models using the latest experience available in the memory buffer $D$, and the Mean Squared Error loss function (Mnih et al. 2015), where $\gamma=0.99$ is a discount factor:
$L_{t}(\theta_{t})={E}_{s,a,r,s^{\prime}}\bigg{[}\bigg{(}R^{t+1}+\gamma\max_{a}Q(s^{t+1},a,\theta_{t})\\ -Q(s^{t},a^{t},\theta_{t})\bigg{)}^{2}\bigg{]}$ (1)
If more than one experience is available (i.e., a player plays more than one dilemma game in that episode), we calculate an average loss across the experiences before updating the network. Each memory buffer $D$ is refreshed before the start of the next episode, so that each player only learns from the latest episode of experience. In updating the network weights with the Adam optimizer (Kingma and Ba 2015), we use a learning rate of $0.001$.
Agents select partners and play the dilemma using an $\epsilon$-greedy policy, acting randomly with probability $\epsilon$ or otherwise playing greedily according to the Q-values learned so far:
We set $\epsilon_{sel}=0.1$ and $\epsilon_{dil}=0.05$. The full detailed algorithm can be found in the Appendix.
Player $M$ can learn according to an extrinsic game reward $R_{M_{extr}}^{t+1}$, which depends directly on the joint actions $a_{M}^{t},a_{O}^{t}$ (as defined in Table 1), or according to an intrinsic moral reward $R_{M_{intr}}^{t+1}$. In the next subsection, we discuss the definition of these intrinsic moral rewards.
Modeling Morality as Intrinsic Reward
@@
Intrinsic rewards associated with pro-social preferences have been used to incentivize the emergence of cooperation in social dilemmas (Hughes et al. 2018; Peysakhovich and Lerer 2018; Jaques et al. 2019). Tennant, Hailes, and Musolesi (2023b) extended this work by defining four Q-learning moral [^1] We use the term moral for agents that are not selfish. However, selfishness itself can be considered as a moral choice, expressed as rational egotism. In the case of social dilemmas, selfishness maps to the concept of rationality. For these agents $R_{M_{intr}}=R_{M_{extr}}$. agents that learn according to various intrinsic rewards $R_{M_{intr}}$, and contrasting these against a traditional Selfish agent which learns to maximize its extrinsic (game) reward $R_{M_{extr}}$ (e.g., Leibo et al. 2017).
@@
In this study, we rely on eight types of moral agents - four pro-social moral learners from Tennant, Hailes, and Musolesi (2023b) (Utilitarian, Deontological, Virtue-Equality, Virtue-Kindness), and four additional anti-social counterparts (anti-Utilitarian, malicious-Deontological, Virtue-Inequality, Virtue-Aggression). The agents are defined as follows (for formal definitions of the reward functions, please refer to Table 2):
@@
  • the Utilitarian (Ut) agent tries to maximize the collective payoff (i.e., total payoff for both players; Bentham 1780);
@@
  • the Deontological (De) agent tries to follow the norm of conditional cooperation (Kant 1785; Fehr and Fischbacher 2004) and gets penalized through the negative reward $-\xi$ for defecting against a cooperator (i.e., an opponent who previously cooperated);
@@
  • the Virtue-Equality (V-Eq) agent tries to maximize equality between the two players’ payoffs $R_{M_{extr}}^{t}$ and $R_{O_{extr}}^{t}$, measured using a two-agent variation of the Gini coefficient (Gini 1912);
@@
  • the Virtue-Kindness (V-Ki) agent receives a reward $\xi$ for acting kindly (i.e., cooperating) against any opponent (Aristotle n.a.);
  • the anti-Utilitarian (aUt) agent tries to minimize the collective payoff;
@@
  • the malicious-Deontological (mDe) agent tries to follow the conditional defection norm and receives a reward $\xi$ for defecting against an cooperator;
@@
  • the Virtue-Inequality (V-In) agent tries to minimize equality between the two players’ payoffs $R_{M_{extr}}^{t}$ and $R_{O_{extr}}^{t}$;
@@
  • the Virtue-Aggression (V-Ag) agent receives a reward $\xi$ for acting aggressively (i.e., defecting) against any opponent.
@@
Table 3: Populations considered in this study. Each population contains eight players of a certain ‘majority’ type and one player of each other type.
Population Compositions
Results